An equation in slope intercept form that passes (5,6) and (3,-4)
Answer by Alan3354(69443)

An equation in slope intercept form that passes (5,6) and (3,-4)
----------------
Find the slope, m
m = diffy/diffx = (-4-6)/(3-5) = -10/-2
m = 5
-----
Use y = mx + b with either point to find b
6 = 5*5 + b
b = -19
-------
y = 5x - 19